Mr. Anders and Ms. Rich each drove home from a business meeting. Mr. Anders traveled east at 100 kilometers per hour and Ms. Rich traveled west at 80 kilometers per hour. In how many hours were they 100 kilometers apart?
Answer by stanbon(75887)

Mr. Anders traveled east at 100 kilometers per hour and Ms. Rich traveled west at 80 kilometers per hour. In how many hours were they 100 kilometers apart?
--------------
Draw the picture.
Anders DATA:
rate = 100 km/h ; distance = x km ; time = d/r = x/100 hrs.
-----------------------
Rich DATA:
rate = 80 km/h ; distance = 100-x km ; time = d/r = (100-x)/80
--------------------------
Equation:
time = time
x/100 = (100-x)/80
80x = 10,000-100x
180x = 10,000
x = 55.56
----------------
time = x/100 = 0.56 hours (time at which they are separated by 100 km)
==============================================
Cheers,
Stan H.